### v1

@@ Appendix A Surface brightness profiles in different bands and length populations
We also split the filament sample into 20–40 Mpc and 40–100 Mpc length populations, we detect a marginal $2.5\sigma$ discrepancy between the two populations (see Fig. 13). The short filaments have a broader stacked emission profile than long filaments, consistent with the predictions in numerical simulations (Galárraga-Espinosa et al. 2021).
@@
Figure 13: Local background subtracted surface brightness profiles of the 20–40 Mpc length (red) and the 40–100 Mpc length (blue) filament samples. The dashed horizontal line denotes the zero level.
Appendix B Impact of cluster masking radius
We investigate the impact of the cluster masking radius on the stacked profile. Instead of the $1.5r_{\lambda}$ masking radii for the optically identified cluster with $\lambda>20$ as described in Sect. 2.1, we use $3r_{\lambda}$ masking radii for a test.
Following Sects. 2.1 and 2.2, we create new 0.3–1.2 keV count and exposure Healpix maps with optically identified clusters masked using $3r_{\lambda}$ masks. Other source masks are the same as those listed in Table 1. We plot the smoothed rate map with $3r_{\lambda}$ masking radii in Fig. 14. Comparing to the count rate map in Fig. 1, the doubled cluster mask size significantly reduced the unmasked footprint.
[…]
Figure 14: Same as Fig. 1, but with $3r_{\lambda}$ mask radii for optically identified clusters with $\lambda>20$.
Figure 15: Comparison between stacked net profile using $1.5r_{\lambda}$ (red) and $3r_{\lambda}$ (blue) cluster mask radii.
@@ Appendix C Mock maps for profile stacking validation
We first phenomenically fit the stacked net surface brightness profile using a King-profile model (see the section on modeling gas properties in the filaments). With the best-fit parameters, we create the mock filament emission rate map based on the projected geometric configurations of the 7817 filaments (top left panel of Fig. 16). We note that the best-fit parameters in this section are empirical, and we do not use them for physical interpretation. The goal is to test the methodology and algorithms used to extract stacked surface brightness profiles. The best-fit profile for injection is plotted as the dotted line in Fig. 5.
We simulate mock background rate maps based on the angular power spectrum of the observed rate map (top right panel of Fig. 16). By comparing the mock filament emission map and the observed sky map, we know that the filament emission contributes $<1\%$ of the total signal. Therefore, we assume that the angular power spectrum of the observed rate map adequately represents that of the background component. We first smooth the rate map using a $1^{\circ}$ full width at half maximum (FWHM) Gaussian kernel to reduce the Poisson noise. Then, we construct a relative amplitude map based on the mean and standard deviation of the smoothed rate map. We compute the angular power spectrum of the relative amplitude map $C_{l,\mathrm{rel}}$ up to a maximum multipole moment $l=3000$ using the function anafast in the HEALPY package. We simulate a mock relative amplitude map using this angular power spectrum and convert it to a mock rate map using the mean and standard deviation of the smoothed rate map. We sum up the mock filament rate map and the mock background rate map to get a mock sky rate map. We obtain a mock sky count map by multiplying the exposure map from observations and applying Poisson randomization to the expected count value in each pixel.
@@
Figure 16: Example of simulated maps in the 0.3–1.2 keV band for validating the imaging stacking method. Top left: mock count rate map for filament emission. Top right: mock foreground/background emission mock map. Bottom left: mock count map of the total emission. Bottom right: observed count map in as a comparison.
@@ Appendix D Spectral models for undetected galactic sources
In Fig. 17, we plot the emission measure distribution of unmasked halos at different redshifts (left panel), the composite spectra of unmasked halos at different redshifts (middle panel), and the averaged spectrum (right panel). In Fig. 18, we plot the spectral models in different redshift bins and stellar mass bins used in the mock catalog, as well as the median spectrum.
Figure 17: Left: Multi-temperature differential emission measure distribution of undetected X-ray halos at different redshifts after applying the selection. Middle: The corresponding composite spectra of each differential emission measure distribution at different redshifts. Right: The synthesis rest-frame spectral model of undetected X-ray halos by taking the filament sample length and redshift distributions into account.

### v2

@@ Appendix A Surface brightness profiles in different bands and length populations
We also split the filament sample into 20–40 Mpc and 40–100 Mpc length populations and we detect a marginal $2.5\sigma$ discrepancy between the two populations (see Fig. 13). The short filaments have a broader stacked emission profile than long filaments, consistent with the predictions in numerical simulations (Galárraga-Espinosa et al. 2021).
@@
Figure 13: Local background-subtracted surface brightness profiles of the 20–40 Mpc length (red) and the 40–100 Mpc length (blue) filament samples. The dashed horizontal line denotes the zero level.
Appendix B Impact of the source masking radius
The choice of source masking radius could lead to a change in the stacked profile. Tanimura et al. (2022) has investigated the impact of different cluster mask radii and point source mask radii on the results. They discovered that the increase in the cluster mask radii from $r_{500}$ to $3r_{500}$ and the increase in the point source mask radii from $30\arcsec$ to $45\arcsec$ do not affect the results (within the uncertainties). In this section, we present our independent check of this systematic uncertainty in our data.
For cluster masks, instead of the $1.5r_{500}$ for the eRASS1 X-ray selected sample and $1.5r_{\lambda}$ for the redMaPPer sample (as described in Sect. 2.1), we increased the masking radii by a factor of two for a test. Following Sects. 2.1 and 2.2, we create new 0.3–1.2 keV count and exposure Healpix maps with clusters masked using $3r_{500}$ and $3r_{\lambda}$ masks. Other source masks are the same as those listed in Table 1. In the left panel of Fig. 14, we plot the stacked net profiles with the doubled cluster masking radii. The profile with doubled mask radii agrees with the original one within uncertainties. The integrated net signal within 10 Mpc transverse distance is 93% of the original integrated signal. We note that the uncertainty in each bin is increased by $13\%$, which is due to the significant reduction of the available footprint (see the comparison between Fig. 1 and Fig. 15).
Similarly, for the eRASS:4 X-ray sources, we tested on the masking radii that are calculated as $S(r_{\mathrm{cut}})=0.05S_{\mathrm{bkg}}$ and plot the comparison in the right panel of Fig. 14. The profiles using two different mask radii are identical within the uncertainties.
Refer to caption Refer to caption
Figure 14: Impact of the source masking radius on the stacked profile. Left: Comparison between stacked net profile using $1.5r_{500}$ (red) and $3r_{500}$ (blue) cluster mask radii. The stacked profile with large cluster masks agrees with the original profile within the uncertainties. Right: Comparison between stacked net profiles using X-ray source masks with $r|_{S=0.1S_{\mathrm{bkg}}}$ (red) and $r|_{S=0.05S_{\mathrm{bkg}}}$ (blue). A more conservative choice of the source mask radius does not change the result within uncertainties.
Figure 15: Same as Fig. 1, but with $3r_{500}$ ($3r_{\lambda}$) mask radii for X-ray selected (optically identified) clusters. The analysis footprint without source masks is significantly reduced.
@@ Appendix C Mock maps for profile stacking validation
First, we phenomenically fit the stacked net surface brightness profile using a King-profile model (see the section on modeling gas properties in the filaments). With the best-fit parameters, we created the mock filament emission rate map based on the projected geometric configurations of the 7817 filaments (top left panel of Fig. 16). We note that the best-fit parameters in this section are empirical and we did not use them for physical interpretation. The goal is to test the methodology and algorithms used to extract stacked surface brightness profiles. The best-fit profile for injection is plotted as the dotted line in Fig. 5.
We simulated mock background rate maps based on the angular power spectrum of the observed rate map (top right panel of Fig. 16). By comparing the mock filament emission map and the observed sky map, we know that the filament emission contributes $<1\%$ of the total signal. Therefore, we assume that the angular power spectrum of the observed rate map adequately represents that of the background component. We first smooth the rate map using a $1^{\circ}$ full width at half maximum (FWHM) Gaussian kernel to reduce the Poisson noise. Then, we construct a relative amplitude map based on the mean and standard deviation of the smoothed rate map. We computed the angular power spectrum of the relative amplitude map $C_{l,\mathrm{rel}}$ up to a maximum multipole moment $l=3000$ using the function anafast in the HEALPY package. We simulated a mock relative amplitude map using this angular power spectrum and convert it to a mock rate map using the mean and standard deviation of the smoothed rate map. We summed up the mock filament rate map and the mock background rate map to get a mock sky rate map. We obtained a mock sky count map by multiplying the exposure map from observations and applying Poisson randomization to the expected count value in each pixel.
@@
Figure 16: Example of simulated maps in the 0.3–1.2 keV band for validating the imaging stacking method. Top left: Mock count rate map for filament emission. Top right: Mock foreground/background emission mock map. Bottom left: Mock count map of the total emission. Bottom right: Observed count map as a comparison.
@@ Appendix D Spectral models for undetected galactic sources
In Fig. 17, we plot the emission measure distribution of unmasked halos at different redshifts (left panel), the composite spectra of unmasked halos at different redshifts (middle panel), and the averaged spectrum (right panel). In Fig. 18, we plot the spectral models in different redshift bins, stellar mass bins used in the mock catalog, and the median spectrum.
Figure 17: Left: Multi-temperature differential emission measure distribution of undetected X-ray halos at different redshifts after applying the selection. Middle: Corresponding composite spectra of each differential emission measure distribution at different redshifts. Right: Synthesis rest-frame spectral model of undetected X-ray halos by taking the filament sample length and redshift distributions into account.